# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ Abstract
In this paper, we study moment and concentration inequalities of the spectral norm for sums of dependent random matrices. We establish novel Rosenthal-Burkholder inequalities for matrix martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. Compared with previous work on matrix concentration inequalities for Markov chains, our results do not require the assumptions of a non-zero absolute spectral gap and bounded matrix functions. Furthermore, our results have leading terms that match the Markov chain central limit theorem, rather than relying on variance proxies. We also give dimension-free versions of the inequalities, which are independent of the ambient dimension $d$ and relies on the effective rank of some matrix instead. This enables the generalization of our results to linear operators in infinite-dimensional Hilbert spaces. Our results have extensive applications in statistics and machine learning; in particular, we obtain improved bounds in covariance estimation and principal component analysis on Markovian data.
@@ 1 Introduction
Concentration inequalities for sums of random matrices (Wigderson and Xiao, 2005; Tropp, 2015) are of great significance for statistics, machine learning, theoretical computer science, signal processing, etc. Common examples include covariance matrix estimation (Chen et al., 2012), principal component analysis (PCA) (Kumar and Sarkar, 2023), and reinforcement learning (Duan and Wainwright, 2022; Wu et al., 2025; Mou, 2025). Although matrix concentration inequalities have been thoroughly studied in the literature (Tropp, 2011, 2012; Mackey et al., 2014; Tropp, 2015; Minsker, 2017; Garg et al., 2018; Klochkov and Zhivotovskiy, 2020; Qiu et al., 2020; Huang and Tropp, 2021; Van Werde and Sanders, 2023; Bandeira et al., 2023; Neeman et al., 2024; Kroshnin and Suvorikova, 2024; Jirak et al., 2024; Wang and Ramdas, 2024; Wu et al., 2025), to the best of our knowledge, the Rosenthal-Burkholder inequality for matrix martingales and matrix Rosenthal inequalities for Markov chains remain absent.
Furthermore, while some existing work (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025) has derived matrix Hoeffding and Bernstein inequalities for Markov chains with a non-zero absolute spectral gap via the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004), these studies suffer from the following limitations. First, for non-reversible Markov chains, geometric ergodicity does not imply a non-zero absolute spectral gap (Kontoyiannis and Meyn, 2012, Theorem 1.4), making these results inapplicable to general non-reversible Markov chains, which occur in reinforcement learning (Duan and Wainwright, 2022) and PCA (Kumar and Sarkar, 2023) problems with Markovian data. Second, these works are confined to uniformly geometrically ergodic Markov chains and bounded matrix functions, thus cannot be extended to unbounded matrix functions. Third, the leading terms of these results do not match the Markov chain central limit theorem (CLT) (Jones, 2004, Corollary 5), relying instead on variance proxies. Finally, due to the use of the multi-matrix Golden-Thompson inequality (Garg et al., 2018, Theorem 1.5), this incurs an additional dependence on the dimension $d$, making it impossible to generalize these results to a dimension-free version. Accordingly, these results cannot be applied to linear operators in infinite-dimensional Hilbert spaces.
@@
The goal of this paper is to bridge the gaps mentioned earlier. Specifically, we first give a Rosenthal-Burkholder inequality (Rosenthal, 1970; Burkholder, 1973; Hitczenko, 1990) for matrix martingales in Theorem 2.1. That is, for an $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, letting martingale differences ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}$ and quadratic variation $\left\langle{{\bm{M}}}\right\rangle_{\infty}=\sum_{k=0}^{\infty}{\mathbb{E}}[{\bm{X}}_{k}^{2}|{\mathscr{F}}_{k-1}]$, one has that
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{R,1}\ \sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big{\|}_{p}+C_{R,2}\ (p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p}.$ (1)
@@
Here, $\|\cdot\|$ is the spectral norm of the matrix, $\|\cdot\|_{p}$ is the $L^{p}$ norm of random variables for any $p\geq 2$, and $C_{R,1},C_{R,2}>0$ are universal constants. This result matches the classic Hilbert space-valued Rosenthal-Burkholder inequality (Pinelis, 1994, Theorem 4.1), with the only difference that an additional $\log d$ term appears. This is an unavoidable price to pay when extending Hilbert-space-valued probabilistic inequalities to matrix versions.
@@
Based on the Rosenthal-Burkholder inequality in Eqn. (1), we further derive matrix Rosenthal inequalities for Markov chains by leveraging the approach of Poisson equations (see Eqn. (12)) (Durmus et al., 2023), which allows us to overcome the limitations mentioned above. In particular, we consider the stationary Markov chain $(Z_{k})_{k\in{\mathbb{N}}}$ in a measurable space $\mathcal{Z}$} with Markov kernel $\mathrm{Q}$} and stationary distribution $\pi$, and a self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$ satisfying $\pi({\bm{F}}):=\int_{{\mathcal{Z}}}{\bm{F}}(z)\pi(dz)=\bm{0}_{d\times d}$. Our goal is to give an upper bound for the spectral norm of the summation ${\bm{S}}_{n}:=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. By the Markov chain CLT (Jones, 2004, Corollary 5), under some regularity conditions, we have $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\pi}({\bm{F}})$, where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\pi}({\bm{F}}):=\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right),$ (2)
@@
with $({\mathrm{Q}}{\bm{F}})(\cdot)=\int_{{\mathcal{Z}}}{\bm{F}}(z){\mathrm{Q}}(\cdot,dz)$. Under the uniform geometric ergodicity assumption (Assumption 2.1) and the geometric $V$-ergodicity assumption with Lyapunov function $V$ (Assumption 2.2), in Theorem 2.2 and Theorem 2.4 we show that for any $p\geq 2$, the following Rosenthal-type inequality holds:
$\frac{1}{n}\left\|{\sup_{k\leq n}\|{\bm{S}}_{k}\|}\right\|_{p}\leq C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right).$ (3)
The first term of the r.h.s. of Eqn.(3) is the leading term, which matches the Markov chain CLT.
@@
By Markov inequality, in Corollary 2.1 and Corollary 2.2 we also translate the Rosenthal inequality (Eqn. (3)) into a Hoeffding and Bernstein high-probability upper-bounds: for any $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
$\frac{1}{n}\sup_{k\leq n}\|{\bm{S}}_{k}\|\leq\frac{16}{3}C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}}{n}}+O\left({\frac{1}{n}}\right),$ (4)
$\frac{1}{n}\sup_{k\leq n}\|{\bm{S}}_{k}\|\leq C_{R,1}\sqrt{\frac{e^{2}\log(\delta^{-1}\vee d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+O\left({\frac{1}{n^{3/4}}}\right),$ (5)
@@
where $t_{\operatorname{mix}}$ is the mixing time and $t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}$ is a variance proxy, which is an upper bound of the long-run variance $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$. Neeman et al. (2024, Theorem 2.6) gave a Bernstein inequality with a leading term relying on the variance proxy $\gamma^{-1}\|\pi({\bm{F}}^{2})\|$ which is also an upper bound of $\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|$, where $\gamma$ is the absolute spectral gap. Hence our results are sharper in the leading term than the previous work on the matrix concentration inequalities for Markov chains (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025) if we consider time-homogeneous matrix functions.
@@
All of the above inequalities (1), (3), (4), (5) can be extended to their respective dimension-free versions, where the ambient dimension $d$ is replaced by the effective rank of some matrix that is an upper bound of the variance term. As a result, our results can be applied to linear operators in infinite-dimensional Hilbert spaces.
We apply our results in covariance estimation, PCA with Markovian data, and obtain improved bounds.
@@
Here and later, $a\wedge b:=\min\{a,b\}$, $a\vee b:=\max\{a,b\}$ for any $a,b\in{\mathbb{R}}$. We also denote ${\mathbb{R}}_{+}:=[0,\infty)$, ${\mathbb{N}}^{*}:=\{1,2,3,\ldots\}$. The symbol $\lesssim$” (resp. $\gtrsim$”) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously.
We denote by ${\bm{I}}_{d}\in{\mathbb{R}}^{d{\times}d}$ the $d\times d$ identity matrix, $\bm{0}_{d_{1}{\times}d_{2}}\in{\mathbb{R}}^{d_{1}{\times}d_{2}}$ the $d_{1}{\times}d_{2}$ zero matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of vector $\bm{u}$}, $\left\|{{\bm{B}}}\right\|$ the spectral norm of matrix $\bm{B}$}, $\bm{B}$$^{*}$ the conjugate transpose of matrix $\bm{B}$}. For any $a\in{\mathbb{R}}$ and a self-adjoint matrix ${\bm{B}}\in{\mathbb{H}}^{d}$ (${\bm{B}}={\bm{B}}^{*}$) with spectral decomposition ${\bm{B}}=\sum_{i=1}^{d}\lambda_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{*}$, we define ${\bm{B}}\wedge a:=\sum_{i=1}^{d}(\lambda_{i}\wedge a){\bm{u}}_{i}{\bm{u}}_{i}^{*}$. For any $\bm{B}_{1},\bm{B}_{2}\in{\mathbb{H}}^{d}$, $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}{-}\bm{B}_{1}$ is positive semi-definite (PSD), and we denote by ${\mathbb{H}}^{d}_{+}$ the set of all $d\times d$ PSD matrices.
We denote by $\mathds{1}$ the indicator function and $\left|{\xi}\right|$ the total variation measure of a signed measure $\xi$. For any $p\geq 1$ and a real-valued random variable $X$, we denote by $\|X\|_{p}=({\mathbb{E}}[|X|^{p}])^{1/p}$ the $L^{p}$ norm of $X$.

### v2

# Matrix Moment and Concentration Inequalities for Martingales and Ergodic Markov Chains with Applications in Statistical Learning

@@ Abstract
In this paper, we study moment and concentration inequalities for the spectral norm of sums of dependent random matrices. We establish novel Rosenthal-Burkholder inequalities for discrete-time matrix local martingales, Burkholder-Davis-Gundy inequality for continuous matrix local martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. Compared with previous work on matrix concentration inequalities for Markov chains, which assume a non-zero absolute $L^{2}$-spectral gap or the stronger $\psi$-mixing condition, our results assume geometric ergodicity, a condition commonly used in statistical applications. Furthermore, our results have leading terms that match the Markov chain central limit theorem, rather than relying on suboptimal variance proxies. We also give dimension-free versions of the inequalities, which are independent of the ambient dimension $d$ and relies on the effective rank instead. This enables the generalization of our results to linear operators in infinite-dimensional Hilbert spaces. Our results have extensive applications in statistics and machine learning; in particular, we obtain improved bounds in covariance estimation and principal component analysis on Markovian data.
@@ 1 Introduction
Moment and concentration inequalities for the spectral norm of sums of random matrices (Wigderson and Xiao, 2005; Tropp, 2015) are of great significance for statistics, machine learning, theoretical computer science, signal processing, etc. Common examples include covariance matrix estimation (Chen et al., 2012), principal component analysis (PCA) (Kumar and Sarkar, 2023), and reinforcement learning (Duan and Wainwright, 2022; Wu et al., 2025; Mou, 2025). Although matrix probability inequalities have been studied extensively in the literature (Tropp, 2011, 2012; Mackey et al., 2014; Tropp, 2015; Paulin et al., 2016; Banna et al., 2016; Minsker, 2017; Tropp, 2018; Garg et al., 2018; Bacry et al., 2018; Klochkov and Zhivotovskiy, 2020; Qiu et al., 2020; Han and Li, 2020; Huang and Tropp, 2021; Cai et al., 2022; Bandeira et al., 2023; Zhivotovskiy, 2024; Neeman et al., 2024; Kroshnin and Suvorikova, 2024; Brailovskaya and van Handel, 2024; Jirak et al., 2025; Wang and Ramdas, 2024; Wu et al., 2025; Maitre, 2025; Van Werde and Sanders, 2025), to the best of our knowledge, the Rosenthal-Burkholder type inequalities for matrix martingales and matrix functions of Markov chains remain absent.
@@
We would bridge the gaps in the current paper. Specifically, we first give a Rosenthal-Burkholder inequality (Rosenthal, 1970; Burkholder, 1973; Hitczenko, 1990) for discrete-time matrix martingales in Theorem 2.1. That is, for an $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted $d$-dimensional self-adjoint matrix martingale $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, letting martingale differences ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}$ and predictable quadratic variation $\left\langle{{\bm{M}}}\right\rangle_{\infty}=\sum_{k=0}^{\infty}{\mathbb{E}}[{\bm{X}}_{k}^{2}|{\mathscr{F}}_{k-1}]$, one has that
[…]
@@
for any $p\geq 2$. Here, $\|\cdot\|$ is the spectral norm of the matrix, $\|\cdot\|_{p}$ is the $L^{p}$ norm of the random variable, and $C_{1},C_{2}>0$ are universal constants. This result matches the classic Hilbert space-valued Rosenthal-Burkholder inequality (Pinelis, 1994, Theorem 4.1), with the only difference that an additional $\log d$ term appears. And the proof is based on the extrapolation method (Theorem 4.1) (Hitczenko, 1990, 1994; Pinelis, 1994) and the techniques in the classic matrix concentration inequalities (Tropp, 2011, 2012, 2015; Kroshnin and Suvorikova, 2024).
As a direct application of our Rosenthal-Burkholder inequality, we further establish a Burkholder-Davis-Gundy inequality for continuous matrix local martingales $({\bm{M}}_{t})_{t\geq 0}$ in Theorem 2.2 by discretization, namely,
$\left\|{\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p},$ (2)
for any $p\geq 2$ and stopping time $\tau$ such that $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}<\infty$.
@@
Regarding Markov chains, based on the Rosenthal-Burkholder inequality in Eqn. (1), we further derive Rosenthal inequalities for sum of time-homogeneous matrix functions of ergodic Markov chains by using the approach of Poisson equations (see Eqn. (14)) (Durmus et al., 2023). In particular, we consider a stationary Markov chain $(Z_{k})_{k\in{\mathbb{N}}}$ in a measurable space $\mathcal{Z}$} with a Markov kernel $\mathrm{Q}$} and a stationary distribution $\mu$, and a square-integrable self-adjoint matrix function of ${\bm{F}}\in L^{2}_{\mu,0}({\mathbb{H}}^{d})$ with $\mu({\bm{F}}^{2}):=\int_{{\mathcal{Z}}}{\bm{F}}^{2}(z)\mu(dz)<\infty$ and $\mu({\bm{F}})=\bm{0}_{d\times d}$. Our goal is to give an upper bound for the spectral norm of the summation ${\bm{S}}_{n}:=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. By the Markov chain central limit theorem (CLT) (Jones, 2004), under some regularity conditions, it holds that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a random matrix with Gaussian entries and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\mu}({\bm{F}})$, where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance defined as
[…]
@@
where $({\mathrm{Q}}{\bm{F}})(\cdot):=\int_{{\mathcal{Z}}}{\bm{F}}(z){\mathrm{Q}}(\cdot,dz)$. Under the uniform geometric ergodicity assumption (Assumption 2.1) and the geometric ergodicity with Lyapunov function $V$ assumption (Assumption 2.2), in Theorem 2.3 and Theorem 2.5 we show that for any $p\geq 2$, the following Rosenthal-type inequality holds:
[…]
where the leading term of the r.h.s. relies on the long-run variance ${\bm{\Sigma}}_{\mu}({\bm{F}})$, matching the Markov chain CLT. Throughout this paper, the $O$-notation captures only the order of $n$ as $n\to\infty$, with other constants omitted.
@@
By the Markov inequality, in Corollary 2.1 and Corollary 2.2 we also translate the Rosenthal inequality (Eqn. (4)) into Hoeffding-type and Bernstein-type high-probability upper-bounds under the uniform geometric ergodicity assumption: for any $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
[…]
@@
where $t_{\operatorname{mix}}$ is the mixing time, and $t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}^{2}$ is a variance proxy which is an upper bound of the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. Neeman et al. (2024, Theorem 2.6) gave a Bernstein inequality with a leading term relying on the variance proxy $\gamma^{-1}\|\mu({\bm{F}}^{2})\|$ which is also an upper bound of $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$, where $\gamma$ is the absolute $L^{2}$-spectral gap. Hence, our results are sharper in the leading term. We apply our results in covariance estimation, PCA with Markovian data, and obtain improved bounds in these applications.
@@
All of the above inequalities (1), (4), (5), (6) can be extended to their respective dimension-free versions, where the ambient dimension $d$ is replaced by the effective rank (see Eqn. (7)) of some matrix that is an upper bound of the variance term. As a result, our results can be applied to linear operators in infinite-dimensional Hilbert spaces.
Finally, we discuss other extensions of these results, which will not be elaborated on in this paper due to space constraints. First, with some efforts, we can prove using the same techniques that the inequality (4) also holds for geometrically ergodic Markov chains with respect to the Wasserstein semimetric (instead of total variation) (Hairer et al., 2011) as in (Durmus et al., 2023). This ergodicity assumption has applications in Markov chain Monte Carlo algorithms (Eberle, 2014; Hairer et al., 2014). Second, based on the Burkholder-Davis-Gundy inequality (2), we can extend the results for Markov chains to continuous Markov processes.
Related Work
Until now, probabilistic inequalities for matrix martingales have been extensively studied. For example, Tropp (2011) proposed a Freedman inequality for matrix martingales; Minsker (2017); Kroshnin and Suvorikova (2024) derived dimension-free matrix Freedman inequalities in different forms, respectively; and Bacry et al. (2018) established a Freedman inequality for continuous-time matrix martingales. Regarding Rosenthal-type moment inequalities, Chen et al. (2012) developed a sharp Rosenthal inequality in the independent case; Jirak et al. (2025) further extend the inequality to its dimension-free version; and Maitre (2025) established a version of Burkholder-Davis-Gundy inequality for continuous matrix martingales. Additionally, in the field of noncommutative probability (Junge and Xu, 2003, 2008; Junge and Zeng, 2013, 2015; Jiao et al., 2025), the probabilistic inequalities for noncommutative martingales proposed therein can also be translated into probabilistic inequalities for the Schatten $p$-norms of matrix martingales. However, the existing results do not imply ours; see the Discussion in Section 2.1 for details.
Regarding probability inequalities for Markov chains, some existing work (Garg et al., 2018; Qiu et al., 2020; Neeman et al., 2024; Wu et al., 2025; Van Werde and Sanders, 2025) has derived matrix concentration inequalities for Markov chains under various ergodicity assumptions. Here we elaborate on the differences between these existing results and ours in detail.
Garg et al. (2018); Qiu et al. (2020); Neeman et al. (2024); Wu et al. (2025) considered Markov chains with a non-zero absolute $L^{2}$-spectral gap (the gap between $1$ and the operator norm of $\mathrm{Q}$} over the subspace $L^{2}_{\mu,0}$), and they used the multi-matrix Golden-Thompson inequality (Sutter et al., 2017; Garg et al., 2018) and the León-Perron argument (León and Perron, 2004) in their proof. However, these studies suffer from the following limitations. First, they assumed a non-zero absolute $L^{2}$-spectral gap $\gamma\in(0,1]$, which is a strong ergodicity condition and implies geometric ergodicity with some Lyapunov function (Assumption 2.2) if we assume the Markov chain is $\psi$-irreducible and aperiodic (Kontoyiannis and Meyn, 2012, Theorem 1.3). This assumption does not hold in many examples, especially for non-reversible Markov chains. For example, Huang and Li (2025, Example A.4) provided a finite-state, non-reversible, irreducible, and aperiodic Markov chain; due to the finiteness of its state space, this Markov chain is uniformly geometrically ergodic (Assumption 2.1), but its absolute $L^{2}$-spectral gap is zero. This strong assumption renders these results inapplicable to general non-reversible Markov chains, which arise in reinforcement learning problems with Markovian data (Duan and Wainwright, 2022). Second, the leading terms of these results rely on suboptimal variance proxies $\gamma^{-1}\|\mu({\bm{F}}^{2})\|$ instead of long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. Finally, the multi-matrix Golden-Thompson inequality (Garg et al., 2018, Theorem 1.5) used in their proof incurs an additional dependence on the dimension $d$, making it impossible to generalize these results to a dimension-free version. Accordingly, these results cannot be easily extended to linear operators in infinite-dimensional Hilbert spaces.
In one recent work, Van Werde and Sanders (2025) considered $\psi$-mixing Markov chains and employed the universality-based approach (Bandeira et al., 2023; Brailovskaya and van Handel, 2024) in their proof. Compared with the classic matrix concentration inequalities derived in our paper, the universality-based concentration inequalities they obtained are sharper in cases involving matrices with specific structures, especially when matrices in the summation are far from being commutative. In their upper bounds, the coefficient of the long-run variance $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$ no longer includes $\log d$. Instead, $\log d$ appears in a new variance term, which is a lower bound of $\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|$. For a detailed introduction of universality-based concentration inequalities, one may refer to the seminal work (Bandeira et al., 2023; Brailovskaya and van Handel, 2024). However, the results in (Van Werde and Sanders, 2025) have the following limitations. On the one hand, $\psi$-mixing is a stronger ergodicity assumption than a non-zero absolute $L^{2}$-spectral gap and uniform geometric ergodicity (Assumption 2.1). Markov chains with infinite states generally fail to satisfy this condition, which limits the scope of application of the results. On the other hand, the approach they employed remains difficult to extend to infinite-dimensional Hilbert spaces.
@@
Here and later, $a\wedge b:=\min\{a,b\}$, $a\vee b:=\max\{a,b\}$ for any $a,b\in{\mathbb{R}}$. We also denote ${\mathbb{R}}_{+}:=[0,\infty)$, ${\mathbb{N}}^{*}:=\{1,2,3,\ldots\}$. The symbol $\lesssim$” (resp. $\gtrsim$”) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously. We denote by $\mathds{1}$ the indicator function and $\left|{\xi}\right|$ the total variation measure of a signed measure $\xi$. For any $p\geq 1$ and a real-valued random variable $X$, we denote by $\|X\|_{p}={\mathbb{E}}^{1/p}[|X|^{p}]$ the $L^{p}$ norm of $X$.
We denote by ${\bm{I}}_{d}\in{\mathbb{R}}^{d{\times}d}$ the $d\times d$ identity matrix, $\bm{0}_{d_{1}{\times}d_{2}}\in{\mathbb{R}}^{d_{1}{\times}d_{2}}$ the $d_{1}{\times}d_{2}$ zero matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of vector $\bm{u}$}, $\left\|{{\bm{B}}}\right\|=\sup_{{\bm{u}}\colon\|{\bm{u}}\|\leq 1}\|{\bm{B}}{\bm{u}}\|$ the spectral norm of matrix $\bm{B}$}, $\bm{B}$$^{*}$ the conjugate transpose of matrix $\bm{B}$}, and ${\bm{B}}]_{ij}$ the $(i,j)$-entry of matrix $\bm{B}$}. For any self-adjoint matrix ${\bm{B}}\in{\mathbb{H}}^{d}$ (${\bm{B}}={\bm{B}}^{*}$) with spectral decomposition ${\bm{B}}=\sum_{i=1}^{d}\lambda_{i}{\bm{u}}_{i}{\bm{u}}_{i}^{*}$, we denote by $\operatorname{Tr}{{\bm{B}}}=\sum_{i=1}^{d}\lambda_{i}$ the trace of matrix $\bm{B}$}, and we define $|{\bm{B}}|:=\sum_{i=1}^{d}|\lambda_{i}|{\bm{u}}_{i}{\bm{u}}_{i}^{*}$ , ${\bm{B}}\wedge a:=\sum_{i=1}^{d}(\lambda_{i}\wedge a){\bm{u}}_{i}{\bm{u}}_{i}^{*}$ for any $a\in{\mathbb{R}}$. For any $\bm{B}_{1},\bm{B}_{2}\in{\mathbb{H}}^{d}$, $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}{-}\bm{B}_{1}$ is positive semi-definite (PSD), and we denote by ${\mathbb{H}}^{d}_{+}$ the set of all $d\times d$ PSD matrices.
Organization.
The remainder of this paper is organized as follows. In Section 2, we present our main results. In Section 3, we apply our results in covariance estimation, PCA problems with Markovian data, and obtain improved error bounds. The proof is outlined in Section 4. In Section 5, we conclude our work. Details of the proof are given in the appendices.